OpenStudy (anonymous):

The latent heat of vaporization of water is 540 calories/gram. How many calories are required to completely vaporize 500 grams of water? 1,040 cal 40 cal 270,000 cal 1.08 cal

OpenStudy (abmon98):

Q=Mass*The latent heat of vaporization Q=540*500=270000 Cal

OpenStudy (anonymous):

Given \[500g\] water

OpenStudy (anonymous):

Conversion factor \[\frac{ 500 \cal}{ 1g}\]

OpenStudy (anonymous):

Solve \[500g \times \frac{ 540cal }{ 1g}= \]